OpenStudy (anonymous):

find the odds of an event if the probability of the event is given to be 1/13

hartnn (hartnn):

The odds in favor of an event or a proposition is defined by the ratio of the probability that the event will happen to the probability that it will not happen. mathematically, if p is the probability, odds = \(\dfrac{p}{1-p}\)

hartnn (hartnn):

here p=1/13, just plug it in!

OpenStudy (anonymous):

So the odds would be \[\frac{ 1 }{ 2 }\]

hartnn (hartnn):

you mean 1/12 ?

hartnn (hartnn):

if you meant 1/12 then you are correct.

OpenStudy (anonymous):

I guess because that is one of the answers, but I got 1/2

OpenStudy (anonymous):

I'm not getting 1/12

OpenStudy (anonymous):

my options are 1:12, 1/13, 12/13, and 12:1

hartnn (hartnn):

ok, \(\huge \dfrac{\dfrac{1}{13}}{1-\dfrac{1}{13}}=\dfrac{1}{13-13\dfrac{1}{13}}=...?\)

hartnn (hartnn):

\(\huge \dfrac{\dfrac{1}{13}}{1-\dfrac{1}{13}}=\dfrac{1}{13-13\dfrac{1}{13}}=\dfrac{1}{13-1}\) so, 1/12 = 1:12 got it ?

OpenStudy (anonymous):

I think so. Thank you. :)

hartnn (hartnn):

welcome ^_^